a. Stephen is walking from his house to his friend Sharon's house. When he is 12 blocks away, he looks at his watch. He looks again when he is 8 blocks away and finds that 6 minutes have passed. Write two ordered pairs for these data in the form (time, blocks).
b. Write a linear equation for these two points.
c. What is the total time it takes Stephen to reach Sharon's house?

Answer by TimothyLamb(4379)

x = time
y = number of blocks remaining to sharon's house
---
(0,12)
(6,8)
---
slope = dy/dx
m = (12-8)/(0-6)
m = -4/6 = -2/3
---
y - 12 = (-2/3)(x - 0)
y = (-2/3)x + 12
---
time to reach sharon's:
0 = (-2/3)x + 12
(2/3)x = 12
x = 12*3/2
x = 18 minutes
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---